OpenStudy (goformit100):

When the mangoes of a basket were counted in twos there was one extra, counted in three were two extra. counted in fours there were three extra, counted in fives there were four extra, counted in sixes there were five extra. But when counted in seven, there were no extra. How many mangoes were there in the basket.

OpenStudy (anonymous):

Since 119/2 remainder=1 119/3 remainder=2 119/4 remainder=3 119/5 remainder=4 119/6 remainder=5 119/7 remainder=0

OpenStudy (anonymous):

Did this years ago

OpenStudy (goformit100):

Thank you Sir.